I need an algebraic expression or formula for solving the following.
" A stamp collection consists of 3, 8, and 15 cent stamps. The number of 8 cent stamps is one less than triple the number of 3 cent stamps. The number of 15 cent stamps is six less than the number of 8 cent stamps. The total value of all the stamps is $2.47. Find the number of 8 cent stamps in the collection".
I was able to solve this using a chart displaying the multiples of 3, 8 and 15 side by side and calculating the sum of the different factors, of course starting by estimating the multiple of 3 first.

Answer by mangopeeler07(462)

x=number of 3 cent stamps
y=number of 8 cent stamps
z=number of 15 cent stamps

3x+8y+15z=247
y=3x-1
z=3x-7

Plug in the values for y and z in the original equation
3x+8(3x-1)+15(3x-7)=247

Distribute
3x+24x-8+45x-105=247

Combine like terms
72x-113=247

Add 113 to both sides
72x=360

Divide both sides by 72
x=5

Those are 3 cent stamps. You want 8 cent stanps, so
y=number of 8 cent stamps
y=3x-1

y=3(5)-1
y=15-1
y=14

There are 14 eight cent stamps in the collection.